When it is not raining, the probability of the Buchanan Bears winning a football game is 2/3. When it is raining the probability of the Bears winning is 3/15. The probability of rain on any day in December is 3/4.
What is the probability of the Bears winning on December 5?


Answer by stanbon(75887)

When it is not raining, the probability of the Buchanan Bears winning a football game is 2/3. When it is raining the probability of the Bears winning is 3/15. The probability of rain on any day in December is 3/4.
What is the probability of the Bears winning on December 5?
-----------------------------------
P(win | not raining) = 2/3
------------------------------
P(win | raining) = 3/15
--------------------------
P(rain) = 3/4
=================================
Question:
P(win) = P(win | rain) + P(win | not raining) = 3/15 + 10/15 = 13/15
==================================
Cheers,
Stan H.